Subscribe to OSU Network Announcements feed
Announcements for OSU's Network and Computer Services
Updated: 28 min 52 sec ago

LS Networks planned maintenance on Midnight to 5 am 11/26/2014

Thu, 11/20/2014 - 9:28am

Primary Contact: Joel Burks,

LSN fiber vendor will be performing an emergency maintenance on November 26th, 2014 00:01 to 02:00 PST in Astoria, OR. LSN fiber vendor will be re-splicing damaged fiber to avoid an potential outage.

Blackboard service restored

Fri, 11/14/2014 - 12:43pm

The Blackboard system is up and running. We are still investigating the root cause of the issue, and how we may prevent it in the future. Thank you for your patience during this time.

-Cody Welch

Blackboard down

Fri, 11/14/2014 - 9:04am

There appears to be a database issue with Blackboard, we are researching it now.

– Lindy Foster Enterprise Computing Oregon State University 541-737-3805

Exchange issues resolved about 1:30PM

Tue, 11/11/2014 - 3:58pm

Exchange service has been restored and appears to be stable at this time.

The issue started today at 11:30AM today and resolved at about 1:30PM. During that time, email clients were connecting to the server but were not receiving new messages.

We are still investigating the root cause of this issue.

-Kirsten Petersen, IT Manager Information Services, Oregon State University p: 541-737-9187, e:

Exchange issues continue – 11/11/2014

Tue, 11/11/2014 - 1:45pm

We are still working to determine the root cause of the intermittent Exchange issues.

We apologize for the continued service disruption.

-Kirsten Petersen, IT Manager Information Services, Oregon State University p: 541-737-9187, e:

Exchange issues 11/11/2014

Tue, 11/11/2014 - 12:20pm

We are seeing intermittent issues with Exchange email and are currently investigating. The cause is not yet known.

-Kirsten Petersen, IT Manager
Information Services, Oregon State University
p: 541-737-9187, e:

11/11 morning maintenance complete

Tue, 11/11/2014 - 9:36am

Primary Contact: Tim Bateman,

The maintenance on the core network from this morning is complete.

Maintenance on OB1

Tue, 11/11/2014 - 4:20am

Primary Contact: Tim Bateman,

Beginning maintenance on OB1. There will be network connectivity loss in the core network this AM during maintenance.

Network slowness

Mon, 11/10/2014 - 4:16pm

Notice of Unplanned Outage

Please be aware that we are currently experiencing high traffic volumes on our network, and some users are experiencing slow downs as a result. We are investigating this, and hope to resolve it as soon as feasible.

If you have questions or concerns about this outage, please contact Network Engineering at net (at) or call 737-HELP.

Nolij is up

Wed, 11/05/2014 - 1:25pm

Nolij service has been restored and we are actively monitoring the application. Thank you for your patience.


Nolij down

Wed, 11/05/2014 - 12:59pm

The Nolij system is currently down. We are looking into the issue and will notify this list when service is restored.


migrating Aruba controller subnet to new core

Tue, 11/04/2014 - 7:30am

Primary Contact: Steven Heitmeyer,

Aruba controller subnet has been successfully moved to the new core. I saw a loss of about 20 pings (= 20 seconds) while routing reconverged. One advantage is that now we have redundant routers. Also we will be able to upgrade to higher data rates in the future.

migrating Aruba controller subnet to new core

Tue, 11/04/2014 - 7:05am

Primary Contact: Steven Heitmeyer,

I am beginning to migrate vlan 87 off of DRT1 (old core) and onto the Nexus 7K routers. There may be a short blip in wireless traffic during the transition.

2014-11-4 06:00-07:00

Mon, 11/03/2014 - 5:11pm

Primary Contact: Richard Giesige,

Network Operations will be conducting an upgrade to the VPN appliances to patch known security vulnerabilities. We do not expect any customer impact, but are sending an outage notice as their is a risk of failover issues during this upgrade period.

2014-11-3 04:00-06:00

Fri, 10/31/2014 - 5:28pm

Primary Contact: Joel Burks,

During the next regular maintenance window, 0400-0600, Monday, 11/03/2014 NERO will be updating firmware on our BGP route reflectors.

We do not expect customer impact, but are sending an outage notice as risk is increased.

2014-11-1 00:00-06:00

Fri, 10/31/2014 - 5:24pm

Primary Contact: Joel Burks,

The circuit provider to NERO’s Ontario hub will perform emergency maintenance between 00:00 and 06:00 on Saturday November 1. The work will replace equipment that is showing signs of failure. Impacted customers include: Malheur ESD, Malheur Ag. Experiment Station, and Malheur Extension office.

Banner Maintenance Saturday November 1 2014

Wed, 10/29/2014 - 12:16pm

We will be upgrading Banner software to address a legislative requirement during our monthly Banner maintenance period this Saturday. Banner will be unavailable November 1 from 12:01AM through 11:59PM.

This outage will affect Banner INB and Banner Self Service (Online Services). Additionally, applications that query or interact directly with Banner may have limited functionality, including Appworx, Cognos, BRM, Cleanaddr, Degree Works (MyDegrees), myOSU, eValS, Nolij, OSCAR, Multiple Assignments, FormFusion, Scholarship Management System, MyVendor, R25 and Banner Workflow.

Jill Swenson Associate Director, Enterprise Computing Services Oregon State University 213 Milne Computer Center Corvallis OR 97331

2014-10-26 00:00-05:00

Fri, 10/24/2014 - 2:35pm

Primary Contact: Joel Burks,

LSN has informed NERO of two upcoming maintenance windows. The first is on Sunday morning, 10/26/2014, between 0000-0500. The second is on Monday, 10/27/2014, between 0000-0500.

During this maintenance we can expect intermittent outages on the backbone circuit between Eugene, Medford, and Klamath Falls. On Sunday we also expect and outage to MetroE connected customers in Eugene.

MetroE connected customers in Eugene: OSU Extension Coos County OSU Extension Linn County OSU Extension Curry County OSU Extension Lincoln County OSU Extension Douglas County

DCA Tools DB migration

Thu, 10/23/2014 - 3:59pm

Primary Contact: Mike Akey,

We will be moving the backend database for the DCA web-based log searching tools during our regular maintenance window from 10pm to 12, Saturday Oct. 25th. During this time, the log searching via tools.nws will be unavailable.

2014-10-23 00:00-06:00

Thu, 10/23/2014 - 9:35am

Primary Contact: Joel Burks,

Level3 has notified NERO of emergency maintenance on damaged fiber between Portland and Seattle. This maintenance will occur tonight, 2014-10-23, between 0000 and 0600.

NERO will monitor the situation and migrate traffic away from the affected circuit as needed.